OpenStudy (anonymous):

If you wanted to eliminate the y variable in the following system, what could be your first step? x-y=6, 4x-3y=12

OpenStudy (texaschic101):

x - y = 6 4x - 3y = 12 The first step would be to multiply the first equation by -3. This would cause the y's to cancel out.

OpenStudy (anonymous):

The first step would be x=6+y. Next step would be 4(6+y)-3y=12.